2,40 Particle 1 is subjected to an acceleration a=-ku, particle 2 is subjected to akt, and particle 3 is subjected to a-ks. All three particles start at the origin s 0 with an initial velocity vo-10 m/s at time t- 0, and the magnitude of k is 0.1 for all three particles (note that the units of k vary from case to e position, velocity, and acceleration ver sus time for each particle over the range 0 sts10 s.
Answer #1

Particle 2

a = -kt

dv/dt = -kt

v = -kt^2/2 + Vo

V = -0.1t^2/2 + 10

Image result for v t graph for decreasing acceleration

particle 1

a = -kv

dv/dt = -kv

dv/v = -kt

logv = -kt + logVo

log(V/Vo) = -kt

V = Voe^(-kt)

= 10e^(-0.1t)

2.5 2.0 1.5 1.0 0.5 0 2 3